An angle is 3 times its complement. Find the two angles.
Answer by macston(5194)

Complementary angles add up to 90 degrees, so:
x+3x=90 degrees
4x=90 degrees Divide each side by 4.
x=22.5 degrees ANSWER 1: The smaller angle is 22.5 degrees.
Larger angle =3 times smaller angle, so
3x=3(22.5 degrees)=67.5 degrees ANSWER 2: The larger angle is 67.5 degrees
And since (22.5 degrees)+(67.5 degrees)=90 degrees, they are complements.